[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Sat Aug 19, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Sat Aug 19 15:12:50 UTC 2017


TAFB Shift Log for Sat Aug 19 2017 11:12:43 GMT-0400 (EDT)

Shift: Day Shift

Pacific Forecaster: M Nelson

Atlantic Forecaster: J Lewitsky

Surface Analyst: M Huffman

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues: The "warehouse" archive stopped working on July 5th. Went  
to look for archived surface analysis maps from last week and noticed that  
none have been archived since July 5th.

Other issues: AC Unit 3 in the server room was alarming but the unit was  
reading 68 degrees compared to the 80 degrees yesterday. All appeared to be  
fine and working despite that alarm today. All Cable boxes in the building  
shut off and powered back on by themselves around 10am. The message on all  
TVs in the WFO and NHC says "This device is currently not activated. Please  
call Comcast".

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170819/d142f155/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list